Find the roots of the equadratic equations by the method of completing the square. Express the roots, if they are not rational, in simplest radical form.
3x2 - 2x - 5 = 0

Answer by fractalier(6550)

From
3x^2 - 2x - 5 = 0
divide by 3 first and get
x^2 - (2/3)x - 5/3 = 0
Now half of 2/3 squared is 1/9, so
x^2 - (2/3)x + 1/9 - 5/3 = 0 + 1/9
x^2 - (2/3)x + 1/9 = 5/3 + 1/9 = 15/9 + 1/9
(x - (1/3))^2 = 16/9
Now take the square root of both sides
x - (1/3) = ± 4/3
x = 1/3 ± 4/3
so that
x = 5/3 or -1